OpenStudy (anonymous):

Don't understand how to do these? If f(x) = x^2 + x, find f(-3). A. -12 B. 3 C. 6

OpenStudy (anonymous):

it just means f(x) = (-3)^2 + x

OpenStudy (shamil98):

f(-3)= (-3)^2 + (-3)

OpenStudy (anonymous):

so its just x + 9

OpenStudy (shamil98):

No it isn't.

OpenStudy (anonymous):

hold up hold up

OpenStudy (anonymous):

its 6

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

@Jaleena Hey, all you have to do is substitute the number in the parentheses (in this case, -3) for ALL of the x's in the original function. Let's go through another, shall we? f(x) = x^2 + 8, find f(12) a) 32 b) 152 c) 76 d) 20 Plug in the value given for x... f(12) = 12^2 + 8 Remember that 12^2 = 12 x 12, not 12 x 2. f(x ) = 144 + 8 f(x) = 152 Let me know if you need more practice! These confused the heck out of me at first, too. And the answer for yours is 6.

OpenStudy (anonymous):

Thank you! I will let you know for sure. @shopgirl1701